Draw the graph of the function f:R→R : f(x)=x²–6x+5. Obtain the coordinates of the points at which the graph meets the x and y axes.
1
Expert's answer
2019-04-12T08:58:32-0400


x intersection:


"y=0, x^2-6x+5=0""x_{1,2}=\frac {6 \pm \sqrt{6^2 - 4 \cdot 5}} {2} = \frac {6 \pm 4} {2}""x_1 =1, x_2 =5""(1,0) and (5,0)"

y intersection:


"x=0, y=0^2 - 5 \cdot 0 +5=5""(0,5)"